OpenStudy (anonymous):

Calculate the discriminant and use it to determine how many real-number roots the equation has. 3x2 – 6x + 3 = 0 A. no real-number roots B. two real-number roots C. three real-number roots D. one real-number root

OpenStudy (anonymous):

@happyrosy

OpenStudy (anonymous):

@Michele_Laino @peachpi

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

you dont have to help me i know the stuff i think i have the answer C

OpenStudy (anonymous):

Alright, same way as before a=3, b=-6, and c=3 (-6)^2-4(3)(3)=36-36=0 Okay, the rule with dsicriminants is: If > 0: There are two real roots If = 0: There is one real root If < 0: There are two complex roots

OpenStudy (anonymous):

The main reason it cannot be C is because the equation is a to the power of two meaning at most, it can only have 2 roots.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D is correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Mhm